OpenStudy (fuzzyball):

Sam opened a money-market account that pays 5% simple interest. He started the account with $7,000 and made no further deposits. When he closed the account, he had earned $700 in interest. How long did he keep his account open? A. 5 years B. 2 years C. 4 years D. 3 years

OpenStudy (fuzzyball):

@FortyTheRapper

OpenStudy (jonathan34):

ok all help you

OpenStudy (fortytherapper):

Let's start off by converting the percent into a decimal. What would 5% be as a decimal

OpenStudy (jonathan34):

#2. $2,145.00 (assuming that interest is simple like the previous question) #3. $12,150.00 #4. $8,000

OpenStudy (fuzzyball):

0.05

OpenStudy (fortytherapper):

Nice, So now we have \[(7000)(.05)= 1 year\] -- $7,000 being the deposit and .05 the interest. Then it should be easy from there

OpenStudy (fuzzyball):

350=1year

OpenStudy (fortytherapper):

Right! Now how can we get 2 years? Same way \[\frac{ 350 }{ 1 }=\frac{ x }{ 2 }\]

OpenStudy (fuzzyball):

350+350=700 so it would be 2 years

OpenStudy (fortytherapper):

Exactly!

OpenStudy (fuzzyball):

thanks so much

OpenStudy (fortytherapper):

Np